In a population of 1000 subjects 640 possess a certain characteristics.A sample of 40 subjects selected from this population has 24 subjects who possess the same characteristics.what are the value of the population and sample proportions.
1
Expert's answer
2021-07-28T07:41:07-0400
"\hat{p}=\dfrac{X}{n}=\dfrac{24}{40}=0.6"

"E[\hat{p}]=p=\dfrac{640}{1000}=0.64"

"\sigma_{\hat{p}}=\sqrt{\dfrac{p(1-p)}{n}}=\sqrt{\dfrac{0.64(1-0.64)}{40}}"

"=0.024\sqrt{10}\approx0.0759"



The value of the population proportion is "p=0.64."


The value of the sample proportion is "\hat{p}=0.6."